Mary, Megan, and Molly sold magazine subscriptions to raise money for playground equipment/ Megan sold 9 more subscriptions than Mary. Molly sold 3 less than twice as many subscribtions as mary. If together they sold a total of 66 magazine subscriptions, how many subscriptions did each girl sell?
Answer by ankor@dixie-net.com(12684)

Mary, Megan, and Molly sold magazine subscriptions to raise money for playground equipment/
:
Write an equation for each statement:
:
Megan sold 9 more subscriptions than Mary.
Me = Ma + 9
:
Molly sold 3 less than twice as many subscriptions as Mary.
Mo = 2Ma - 3
:
If together they sold a total of 66 magazine subscriptions,
Ma + Me + Mo = 66
:
how many subscriptions did each girl sell?
:
In the above equation: Replace Me with (Ma+9); and replace Mo with (2Ma-3)
Ma + (Ma+9) + (2Ma-3) = 66
Ma + Ma + 2Ma + 9 - 3 = 66
4Ma + 6 = 66
4Ma = 66 - 6
4Ma = 60
Ma = 60/4
Ma = 15 Sold by Mary
:
I'll let you find Mo and Me